OpenStudy (anonymous):

Find the volume that is restricted by paraboloid z = 2x^2 +2y^2 -8 and xy-plane using polar coordinates.

ganeshie8 (ganeshie8):

start by changing to polar

ganeshie8 (ganeshie8):

V = \(\large \int \int_R z ~dx dy\) u need to convert z, differentials, and bounds to polar..

ganeshie8 (ganeshie8):

\(x = r\cos \theta\) \(y = r\sin \theta\) convert z to polar, wat do u get ?

OpenStudy (anonymous):

\[2r ^{2} -8\]

ganeshie8 (ganeshie8):

yes, next think a bit how the projection of paraboloid look in xy-axis

ganeshie8 (ganeshie8):